OpenStudy (anonymous):

A parachutist's rate during a free fall reaches 225 kilometers per hour. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 15 seconds of free fall? Do not round your answers. RATE: DISTANCE FALLEN IN 15 SECONDS

OpenStudy (anonymous):

937.5

OpenStudy (anonymous):

HOW BOUT DISTANCE?

OpenStudy (anonymous):

@Xander22

OpenStudy (johnweldon1993):

\[\large 225 \frac{\cancel{km}}{\cancel{hr}} \times \frac{1 \cancel{hr}}{3600s} \times \frac{1000m}{1\cancel{km}} = 62.5 \frac{m}{s}\]

OpenStudy (johnweldon1993):

To find out the distance, just realize if the parachute falls 62.5 meters every second...after 15 seconds how many meters has it fallen?